OpenStudy (anonymous):

the sales tax on the purchase of a futon is $33.39. if the tax rate is 7% find the purchase price of the futon

OpenStudy (turingtest):

33.39+(7/100)(33.39)=$35.73

OpenStudy (anonymous):

it says that is incorrect

OpenStudy (turingtest):

oh my bad, I misread, let me try again...

OpenStudy (turingtest):

0.07x=33.39 x=477 (the price of the futon before tax) total price is the futon price plus tax: 477+33.39=$510.39